Solve the triangle.

A = 21°, C = 105 °, c = 5


1
Expert's answer
2022-05-17T12:38:26-0400

B=1800-1050-210=540

a/sinA=b/sinB=c/sinC

a/sin210=b/sin540=5/sin1050

a=5sin210/sin1050=5*0.3589/0.9659=1.85

b=5sin540/sin1050=5*0.809/0.9659=4.18